### v1

# On Signed Network Coordination Games

Abstract
We study binary-action pairwise-separable network games that encompass both coordinating and anti-coordinating behaviors. Our model is grounded in an underlying directed signed graph, where each link is associated with a weight that describes the strenght and nature of the interaction. The utility for each agent is an aggregation of pairwise terms determined by the weights of the signed graph in addition to an individual bias term. We consider a scenario that assumes the presence of a prominent ’cohesive’ subset of players, who are either connected exclusively by positive weights, or forms a structurally balanced subset that can be bipartitioned into two adversarial subcommunities with positive intra-community and negative inter-community edges. Given the properties of the game restricted to the remaining players, our results guarantee the existence of Nash equilibria characterized by a consensus or, respectively, a polarization within the first group, as well as their stability under best response transitions. Our results can be interpreted as robustness results, building on the super-modular properties of coordination games and on a novel use of the concept of graph cohesiveness.
Index terms: Coordination games, anti-coordination games, network games, signed graphs, network robustness, best response dynamics, robust stability.
I Introduction
A key feature of many socio-technical systems is the heterogeneity among the behaviors of the agents in the network and among their interactions and mutual influences. Such heterogeneities pose significant challenges in various fields including traffic and routing games, epidemics models.
In this paper, we focus on heterogeneous interactions within networks of agents engaged in strategic games with binary actions. Our model encompasses two prominent families of network games [3, 4, 5, 6]: network coordination games [7, 8, 9, 10, 11] and network anti-coordination games [12, 13, 14, 15, 16], both of which have a variety of applications in economics, social sciences, and biology. In their simplest version, the utility of a player in a network coordination (anti-coordination) game is an affine increasing (decreasing) function of the number of her neighbors in the network playing the same action. Despite their apparent similarity, both fundamental properties and applications of network coordination and network anti-coordination games are quite different.
Network coordination games model the so called strategic complements effects, that is when the choice of a certain action by one player makes it more appealing for other players to play the same action. They are used to model social network features like the adoption of beliefs or behavioral attitudes, or economic ones such as the spread of a new technology. Mathematically, they belong to the broader class of super-modular games [17, 18, 19, 20]. As a consequence, Nash equilibria always exist and one special instance of them are the consensus strategy profiles, namely, those where all individuals are playing the same action. Moreover, asynchronous best response dynamics globally reach the set of Nash equilibria in network coordination games. [21]
In contrast, network anti-coordination games are representative of another class of games exhibiting the so called strategic substitutes effect. In this case, the choice of a certain action by one player makes it more appealing for the other players to play the opposite action. They provide a natural model in situations where players are competing for resources that can become congested or in models where players can provide a public good, buy snob goods, or, in general, when there are gains from differentiation. In contrast to network coordination games, existence of Nash equilibria for anti-coordination games is guaranteed in special cases strongly dependent on the network structure. [14] Moreover, even when equilibria exist, asynchronous best response dynamics may have limit cycles.
Network games comprising both coordinating and anti-coordinating interactions have been recently proposed to model the presence of anti-conformist behaviors in a social community, accounting for some form of heterogeneity [22, 23, 24, 25] . More generally, games exhibiting both strategic complements and substitutes have been recently proposed in the economic literature [26] to model heterogeneous interactions, e.g., markets with coexistence of both Cournot and Bertrand type firms. Such mixed games may fail to possess Nash equilibria. A fundamental example is the matching pennies game, which is a two-player game with one coordinating and one anti-coordinating player. In contrast with the matching pennies game, one can imagine that in a scenario where most of the players are coordinating and form a ’well’ connected subset, the presence of few anti-coordinating players should not prevent the coordinating players to reach a consensus and possibly the whole system to reach a Nash equilibrium. This is one main motivation to our work.
Another application is when a subset of the graph is structurally balanced [27, 28]. A structurally balanced signed graph is one where vertices can be split into two subsets so that intra links on every set have positive weight, while links connecting the two groups have negative weight. Such property envisages a polarization in the system’s equilibrium. In this case we want to determine whether such polarization is still reached even if the entire graph is not structurally balanced.
Network models with signed weights have appeared in many other different fields: to model the presence of antagonistic interactions in social networks [27, 28, 29, 30], inhibitory signals in genetics [31, 32] and neural networks [33], antiferromagnetic bonds in spin glasses [34]. The popular Linear Threshold Model (LTM) originally introduced in [35] for non negative graphs has been recently proposed in the context of signed graphs in [36, 37]. This dynamical system is strictly related to the best response dynamics for network ccordination games. Literature on LTM on signed graphs has focused on showing the differences with respect to the nonnegative case: lack of cascades, dependence on the activation pattern, polarization. Somewhat analogous are the studies of the linear averaging dynamics on signed graphs [38, 39]. Therein, the important concept of structurally balanced graph is used to determine the structure of the steady state dynamics.
In this paper, we consider a finite set of agents whose network of interactions is modelled as a directed signed graph. Given two agents, it is possible that a link exists in just one direction and, even when both links are present, they may have a different weight and possibly weights with opposite signs. This means that agent $i$ may tend to coordinate with an agent $j$, while agent $j$ tends to anti-coordinate with $i$. The utility of agent $i$ is an aggregation of a family of pairwise terms one for each of the out-neighbors of $i$ plus an individual bias term. Each pairwise term can be of coordination or anti-coordination type and is modulated by the corresponding graph weight. We call such games signed network coordination games. The focus of this paper is on the existence of pure strategy Nash equilibria in signed-network coordination (SNC) games and the analysis of their stability with respect to best response transitions.
Our work builds on two fundamental concepts. The first one is a novel use of the notion of cohesiveness originally proposed in the pioneering work [9] to describe the pure Nash equilibria of network coordination games. The second is the super-modularity property [17, 18, 19, 20] of network coordination games, particularly the robustness results recently appeared in [40]. The general setup of our results is that of a SNC game where the set of agents $\mathcal{V}$ is split into two subsets $\mathcal{V}=\mathcal{R}\cup\mathcal{S}$. Agents in $\mathcal{R}$ are assumed to be intra themselves coordinating and form a sufficiently ’cohesive’ subset. Alternatively, they are assumed to form a ’cohesive’ structurally balanced subset. In this latter case, a transformation of the strategy profile set allows to obtain a coordinating subset. If agents in $\mathcal{S}$ possess an equilibrium conditioned to any possible value of the actions taken by the agents in $\mathcal{R}$, then a Nash equilibrium exists that is a consensus or a polarization (respectively) on the agents in $\mathcal{R}$. This is the content of our first result, Theorem 1.
Conditions for the convergence of the best response dynamics are characterized in terms of a novel notion of indecomposability, related to the uniform non-cohesiveness property used in [9], and make a fundamental use of robust properties of super-modular games [40]. Our main result, Theorem 2, formalizes these ideas.
Weaker preliminary results have appeared in [1] and [2]. In [1], the underlying graph of interactions was a complete one. In [2], each agent was engaged in interactions of only one type, either coordinating or anti-coordinating. Moreover, the graph restricted to each of the two subgroups was assumed to be undirected.
We conclude this section by presenting a brief outline of this paper. The remainder of this section is devoted to the introduction of some basic notational conventions to be followed throughout the paper. In Section II, we formally introduce SNC game and we present few examples illustrating the problem we want to consider. In Section III, we derive some preliminary results for three classes of signed graphs: undirected, unsigned and structurally balanced. In Section IV, we present our main results, in particular Theorems 1 and 2, and show their applicability. Finally, Section V presents some final remarks.
Notation
Let $\mathds{R}$ and $\mathbb{R}_{+}$ denote the set of real numbers and non-negative real numbers, respectively. For a finite set $\mathcal{U}$, $\mathbb{R}^{\mathcal{V}}$ denotes the space of real column vectors $x$, whose entries $x_{i}$ are indexed by the elements $i$ in $\mathcal{V}$. For a vector $x$ in $\mathbb{R}^{\mathcal{V}}$ and a subset $\mathcal{A}\subseteq\mathcal{V}$, we use the notation $x_{\mathcal{A}}$ in $\mathbb{R}^{\mathcal{A}}$ for the restricted vector with entries $(x_{\mathcal{A}})_{i}=x_{i}$ for every $i$ in $\mathcal{A}$. The all-$1$ vector, whose size may be deduced from the context, will be denoted by $\bm{1}$. For a vector $x$ in $\mathbb{R}^{\mathcal{V}}$ and some $i$ in $\mathcal{V}$, we write $x_{-i}=x_{\mathcal{V}\setminus\{i\}}$ for the vector in $\mathbb{R}^{\mathcal{V}\setminus\{i\}}$ obtained from $i$ by removing its $i$-th entry.
Similarly, for two finite sets $\mathcal{U}$ and $\mathcal{V}$, $\mathbb{R}^{\mathcal{U}\times\mathcal{V}}$ denotes the space of real matrices $M$ whose entries $M_{ij}$ are indexed by the pairs $(i,j)$ in $\mathcal{U}\times\mathcal{V}$. For a matrix $M$ in $\mathbb{R}^{\mathcal{U}\times\mathcal{V}}$ and two subsets $\mathcal{A}\subseteq\mathcal{U}$ and $\mathcal{B}\subseteq\mathcal{V}$, we use the notation $M_{\mathcal{A}\mathcal{B}}$ in $\mathbb{R}^{\mathcal{A}\times\mathcal{B}}$ for the restricted matrix, with entries $(M_{\mathcal{A}\mathcal{B}})_{ij}$ for every $i$ in $\mathcal{A}$ and $j$ in $\mathcal{B}$. For a vector $d$ in $\mathbb{R}^{\mathcal{A}}$, $[d]$stands for the diagonal matrix with diagonal coinciding with $d$, i.e., $[d]\in\mathbb{R}^{\mathcal{A}\times\mathcal{A}}$ is such that $[d]_{ii}=d_{i}$ and $[d]_{ij}=0$ for every $i$ in $\mathcal{A}$ and $j$ in $\mathcal{A}\setminus\{j\}$.
II Model
In this section, we introduce the model in its general form and present the main issues addressed in the rest of the paper.
II-A Signed Network Coordination Games
We model networks as finite directed weighted signed graphs $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$, with non-empty set of nodes $\mathcal{V}$, set of directed links $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, and weight matrix $W$ in $\mathbb{R}^{\mathcal{V}\times\mathcal{V}}$ such that $W_{ij}\neq 0$ if and only if $(i,j)$ is a link in $\mathcal{E}$. We do not allow for the presence of self-loops, equivalently, we assume that the weight matrix $W$ has zero diagonal. Throughout, finite directed weighted signed graphs without self-loops will be simply referred to as networks. [^1] Notice that we do not assume in general that $W_{ij}$ and $W_{ji}$ have the same sign (a property referred to as “digon sign-symmetry” in [39].)
We shall call a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ undirected when the weight matrix $W=W^{\prime}$ is symmetric, so that in particular $(j,i)$ is a link in $\mathcal{E}$ whenever $(i,j)$ is, and in this case they have the same weight $W_{ij}=W_{ji}$. We refer to a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ as unsigned when the weight matrix $W$ is non-negative, i.e., all links $(i,j)$ in $\mathcal{E}$ have positive weight $W_{ij}>0$. For a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ and a subset of nodes $\mathcal{U}\subseteq\mathcal{V}$, we consider the subnetwork $\mathcal{G}_{\mathcal{U}}=(\mathcal{U},\mathcal{E}_{\mathcal{U}},W_{\mathcal{U}\mathcal{U}})$ where $\mathcal{E}_{\mathcal{U}}=\mathcal{E}\cap(\mathcal{U}\times\mathcal{U})$ is the subset of links with both tail node and head node in $\mathcal{U}$, while $W_{\mathcal{U}\mathcal{U}}$ is the sub-matrix of $W$ obtained by restricting its row and column sets to $\mathcal{U}$.
Example 1.
Figure 1 illustrates a network with weight values reported next to the corresponding links. Notice that, for =$\mathcal{R}=\{1,\dots,8\}$ and ~=$\tilde{\mathcal{R}}=\{1,\dots,6\}$, the subnetworks $\mathcal{G}_{\mathcal{R}}$ and $\mathcal{G}_{\tilde{\mathcal{R}}}$ are both unsigned, whereas for $\mathcal{S}=\mathcal{V}\setminus\mathcal{R}=\{9,\dots,13\}$ and $\tilde{\mathcal{S}}=\mathcal{V}\setminus\tilde{\mathcal{R}}=\{7,\dots,13\}$ the subnetworks $\mathcal{G}_{\mathcal{S}}$ and $\mathcal{G}_{\tilde{\mathcal{S}}}$ are both undirected.
134562789101112132231-2-7-2 3212-14-413-2-5-1-1-13
Figure 1: A network with 13 nodes. Weights are represented by the values on the links. Nodes in the set =$\mathcal{R}=\{1,\dots,8\}$ are colored in gray, while nodes in the complement set $\mathcal{S}=\{9,\dots,13\}$ are white.
Definition 1.
The signed network coordination (SNC) game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$ in $\mathbb{R}^{\mathcal{V}}$ is the strategic game with player set $\mathcal{V}$, whereby every player $i$ in $\mathcal{V}$ has action set $\mathcal{A}=\{\pm 1\}$ and utility function $u_{i}:\mathcal{X}\rightarrow\mathbb{R}$ given by
$u_{i}(x)=u_{i}(x_{i},x_{-i})=h_{i}x_{i}+x_{i}\sum_{j\in\mathcal{V}}W_{ij}x_{j}\,,$ (1)
for every strategy profile $x$, where $\mathcal{X}=\mathcal{A}^{\mathcal{V}}$ denotes the set of strategy profiles of all players and $x_{-i}$ in $\mathcal{X}_{-i}=\mathcal{A}^{\mathcal{V}\setminus\{i\}}$ stands for the strategy profile of all players except for player $i$.
We now provide an interpretation of the SNC games introduced above. The utility function (1) is the sum of two terms. The first addend in the right-hand side of (1) is a standalone term $h_{i}x_{i}$ (i.e., a term that depends only on the action of player $i$) that models the bias of player $i$ towards one of the two possible actions depending on the sign of the preference term $h_{i}$: if $h_{i}>0$ (respectively, $h_{i}<0$), then player $i$ has a bias towards action $+1$ ($-1$), in that, in the absence of the network (i.e., if $W=0$) this action would provide a utility that is $2h_{i}$ larger that its alternative $-1$ ($+1$); if $h_{i}=0$, then no bias is present in that both actions $+1$ and $-1$ would reward player $i$ with the same utility in the absence of the network. The second addend in the right-hand side of (1) is the aggregate of pairwise interaction terms $x_{i}W_{ij}x_{j}$, each of which can be interpreted as the utility of player $i$ in a two-player game with player $j$. Notice that, while the sum index $j$ in the right-hand side of (1) formally runs over all players $j$ in $\mathcal{V}$, the sum is effectively only over players $j$ such that $W_{ij}\neq 0$, i.e., out-neighbors of player $i$ in the network $\mathcal{G}$. Observe that a SNC game with binary actions is an instance of a polymatrix or pairwise-separable game. [41, 42].
We say that a player $i$ in $\mathcal{V}$ has a coordinating interaction with another player $j$ when $W_{ij}>0$, meaning that the utility $u_{i}(x)$ of player $i$ increases when player $j$ plays the same action as $i$, i..e, when $x_{i}x_{j}=1$. Conversely, when $W_{ij}<0$, player $i$ is said to have an anti-coordinating interaction with player $j$ in $\mathcal{V}$, namely its utility decreases when player $j$ is playing the same action. In the special case when the network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ is unsigned, namely when $W_{ij}\geq 0$ for every $i$ and $j$ in $\mathcal{V}$, the SNC game with binary actions is known as a network coordination game. On the other hand, when $W_{ij}\leq 0$ for every $i$ and $j$ in $\mathcal{V}$, the SNC game is known as a network anti-coordination game.
Observe that our general model of SNC game accounts for asymmetric behaviors: while it is possible that two nodes $i$ and $j$ have the same type of interaction one against the other one, it can happen that a player $i$ has a coordinating interaction with another player $j$ (i.e., $W_{ij}>0$) while $j$ has an anti-coordinating interaction with $i$ or is not influenced by player $i$ at all (i.e., $W_{ji}=0$). This makes the analysis of such games particularly challenging. Notice that our framework encompasses and generalizes the special cases studied in our previous works [1] and [2] where, for every player $i$ in $\mathcal{V}$, either $W_{ij}\geq 0$ for every $j$ in $\mathcal{V}$, or $W_{ij}\leq 0$ for every $j$ in $\mathcal{V}$.
II-B Best Responses and Nash Equilibria
The best response (BR) correspondence for a player $i$ in $\mathcal{V}$ is the set of its optimal actions given the strategy profile or the other players. Formally,
$\mathcal{B}_{i}(x_{-i})=\operatorname*{argmax}_{x_{i}\in\mathcal{A}}u_{i}(x_{i},x_{-i})\,.$
A (pure strategy) Nash equilibrium is a strategy profile $x^{*}$ in $\mathcal{X}$ such that
$x^{*}_{i}\in\mathcal{B}_{i}(x^{*}_{-i})\,,\qquad\forall i\in\mathcal{V}\,.$
The set of Nash equilibria of a game will be denoted as $\mathcal{N}$. A Nash equilibrium $x^{*}$ in $\mathcal{N}$ is called strict if $\mathcal{B}_{i}(x^{*}_{-i})=\{x^{*}_{i}\}$ for every $i$ in $\mathcal{V}$. The set of strict Nash equilibria of a game will be denoted as $\mathcal{N}^{*}$.
The purpose of this work is to investigate the existence of Nash equilibria for SNC games and study their stability properties under best response transformations. Formally, consider two strategy profiles $x$ and $y$ in $\mathcal{X}$. For $l\geq 0$, a length-$l$ best response path (BR-path) from $x$ to $y$ is an $(l+1)$-tuple of strategy profiles in $\mathcal{X}^{l+1}$, denoted with $(x^{(0)},x^{(1)},\ldots x^{(l)})$, such that
  • $x^{(0)}=x$, and $x^{(l)}=y$;
  • for every $k=1,2,\ldots,l$, there exists a player $i_{k}$ in $\mathcal{V}$ such that
    $x^{(k)}_{-i_{k}}=x^{(k-1)}_{-i_{k}}\,,\qquad x_{i_{k}}^{(k)}\in\mathcal{B}_{i_{k}}(x_{-i_{k}}^{(k-1)})\setminus\{x^{(k-1)}_{i_{k}}\}\,.$ (2)
A BR-path is thus a sequence of unilateral modifications of the strategy profile where one agent at a time changes its action to any of its best responses to the current strategy profile of the rest of the population. Notice that, with the above definition, for every strategy profile $x$ in $\mathcal{X}$, there always exists a length-00 BR-path $(x)$ from $x$ to itself.
We shall refer to a subset of strategy profiles $\mathcal{X}^{*}\subseteq\mathcal{X}$ as:
  • BR-reachable from strategy $x$ in $\mathcal{X}$ if there exists a BR-path from $x$ to some strategy profile $y$ in $\mathcal{X}^{*}$;
  • globally BR-reachable if it is BR-reachable from every strategy profile $x$ in $\mathcal{X}$;
  • BR-invariant if there are no BR-paths from any $y$ in $\mathcal{X}^{*}$ to any $z$ in $\mathcal{X}\setminus\mathcal{X}^{*}$;
  • globally BR-stable if it is globally BR-reachable and BR-invariant.
Notice that a BR-reachable (hence, in particular a globally BR-reachable or globally BR-stable) set of strategy profiles $\mathcal{X}^{*}$ is necessarily non-empty. Observe that a singleton $\{x^{*}\}$ is BR-invariant if and only if $x^{*}$ is a strict Nash equilibrium. A globally BR-stable set is one from which it is impossible to exit through unilateral best response modifications and that can be reached from every initial strategy profile through a finite number of such modifications.
Remark 1.
The notions of BR-reachability, BR-invariance, and BR-stability are particularly relevant for the study of so-called asynchronous best response dynamics [43]. These are a class of asynchronous dynamics modeled as discrete-time Markov chains with finite state space coinciding with the strategy profile set $\mathcal{X}$, whereby, at every time step $t=0,1,2,\ldots$, conditioned on the current configuration $X(t)=x$, one player $i$ is randomly selected from the player set $\mathcal{V}$ according to a distribution that assigns positive probability to all players, and she updates her action to a value $X_{i}(t+1)$ chosen uniformly from her best response set $\mathcal{B}_{i}(x_{-i})$. In fact, observe that, on the one hand, a subset of strategy profiles $\mathcal{X}^{*}\subseteq\mathcal{X}$ is BR-invariant if and only if it is a trapping set for the asynchronous best response dynamics, i.e., if $X(s)\in\mathcal{X}^{*}$ for some $s\geq 0$ implies that $X(t)\in\mathcal{X}^{*}$ for every $t\geq s$. On the other hand, $\mathcal{X}^{*}$ is globally BR-stable if and only if, for every probability distribution of the initial state $X(0)$ of the asynchronous best response dynamics, there exists a random time $T$ that is finite with probability $1$ and is such that $X(t)\in\mathcal{X}^{*}$ for every $t\geq T$.
Remark 2.
In [40], the notion of improvement path (I-path) was introduced along with the related notions of I-reachability, I-invariance, and I-stability. Specifically, a length-$l$ I-path is a $(l+1)$-tuple of strategy profiles $(x^{(0)},x^{(1)},\ldots x^{(l)})$ such that for every $k=1,2,\ldots,l$ there exists a player $i_{k}$ in $\mathcal{V}$ such that
$x^{(k)}_{-i_{k}}=x^{(k-1)}_{-i_{k}}\,,\qquad u_{i_{k}}(x^{(k)})>u_{i_{k}}(x^{(k-1)})\,.$ (3)
Notice that, if the action space of every player is binary, as is our case, then condition (3) implies condition (2), so that every I-path is also a BR-path, but not necessarily vice versa. As a consequence, every (globally) BR-reachable set of strategy profiles is (globally) BR-reachable. Conversely, every BR-invariant set of strategy profiles is I-invariant, but not vice versa: e.g., every non-strict Nash equilibrium is I-invariant but not BR-invariant. On the other hand, neither I-stability implies BR-stability nor vice versa: e.g., for the SNC game of Example 4 with $\alpha=0$, the set of Nash equilibria $\mathcal{N}$ is globally I-stable but not BR-stable.
The following two examples show how Nash equilibria may not exist for SNC games.
121-1(a)
123-1-1-1(b)
Figure 2: Two signed graphs representing: (a) the discoordination game and (b) a directed anti-coordination game.
Example 2.
The SNC game on the network (a) in Figure 2 with external field $h=0$ reduces to a two-player game with binary action set $\mathcal{A}=\{\pm 1\}$ and utilities $u_{1}(x_{1},x_{2})=x_{1}x_{2}$ and $u_{2}(x_{1},x_{2})=-x_{1}x_{2}$, which is commonly referred to as the discoordination game [44].
Example 3.
The SNC game with binary actions on the network (b) in Figure 2 with external field $h=0$ is an anti-coordination game. For $x^{*}$ to be a Nash equilibrium, one should have $x_{1}^{*}=-x_{2}^{*}=x_{3}^{*}=-x_{1}^{*}$, which is impossible in {±$\{\pm 1\}^{3}$. Hence, this game does not admit any Nash equilibria.
Even for SNC games that do admit Nash equilibria, these may fail to be globally BR-reachable or BR-invariant, as illustrated in the following two examples.
Example 4.
Consider a SNC game on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with two nodes $\mathcal{V}=\{1,2\}$, a single link =$\mathcal{E}=\{(1,2)\}$, and weight matrix $W=(1,0)(0,1)^{\prime}\,,$ and external field $h=(0,\alpha)$, where $\alpha$ is a scalar parameter. Then, the set of Nash equilibria is
$\mathcal{N}=\left\{\begin{array}[]{lcl}\{-\bm{1}\}&\text{ if }&\alpha<0\,,\\ \{\pm\bm{1}\}&\text{ if }&\alpha=0\,,\\ \{\bm{1}\}&\text{ if }&\alpha>0\,.\end{array}\right.$
It easily verifiable that $\mathcal{N}$ is globally BR-reachable for every $\alpha$. In fact, for $\alpha\neq 0$ then the unique Nash equilibrium is strict, so that $\mathcal{N}$ is BR-invariant, hence BR-stable. On the other hand, for $\alpha=0$, there are two Nash equilibria neither of which is strict: in this case, neither $\mathcal{N}$ nor any of it subsets are BR-reachable, and there are no BR-stable sets.
Example 5.
812345671111-1-11-11111
Figure 3: A signed graph with $8$ nodes and a coordinating set made of $7$ players (in gray).
Consider the signed-network related to the graph in Figure 3 with $h=0$. Observe that the set =$\mathcal{R}=\{1,\dots,7\}$ is a coordinating set. A direct check shows that the set of Nash equilibria is $\mathcal{N}=\{\pm x^{*}\}$, where $x^{*}$ is the strategy profile with $x^{*}_{\mathcal{R}}=\bm{1}$ and $x^{*}_{8}=-1$. However, $\mathcal{N}$ is not globally BR-reachable, as the subset of strategy profiles
$\mathcal{X}_{\mathcal{R}}=\{y\in\mathcal{X}:\,y_{1}=y_{2}=y_{3}=-y_{5}=-y_{6}=-y_{7}\}\,,$
is BR-invariant.
III Preliminary results
In this subsection, we shall first present some preliminary results on two special classes of SNC games: SNC games on undirected networks and network coordination games (i.e., SNC games on unsigned networks). We will rely on standard results on super-modular and potential games, respectively, to show that the set of Nash equilibria $\mathcal{N}$ of both these classes of games is non-empty and globally BR-reachable, and, for undirected SNC games, that $\mathcal{N}$ contains a non-empty globally BR-stable subset. We will then present a number of examples of SNC games outside these two classes that illustrate the specific challenge in the study of the existence and stability of Nash equilibria of general SNC games.
III-A SNC games on undirected networks
A strategic game with strategy profile space $\mathcal{X}$ is called exact potential if there exists a potential function $\Phi:\mathcal{X}\to\mathbb{R}$ such that, for every player $i$ in $\mathcal{V}$ and strategy profiles $x$ and $y$ in $\mathcal{X}$,
$x_{-i}=y_{-i}\quad\Longrightarrow\quad u_{i}(y)-u_{i}(x)=\Phi(y)-\Phi(x)\,,$ (4)
It is well known [45] that the set of Nash equilibria $\mathcal{N}$ of every finite exact potential game is always nonempty: in particular, every global maximum point $x^{*}$ of the potential function $\Phi(x)$ is a Nash equilibrium. The following result states that a SNC game on a network $\mathcal{G}$ is an exact potential game if and only if $\mathcal{G}$ is undirected and that, in this case, its set of Nash equilibria contains a nonempty globally BR-stable subset.
Proposition 1.
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$. Then,
  1. (i)
    the game is exact potential if and only if $\mathcal{G}$ is undirected.
Moreover, if $\mathcal{G}$ is undirected, then:
  1. (ii)
    $\Phi:\mathcal{X}\to\mathbb{R}$ is an exact potential function for the SNC game if and only if there exists a constant $C$ in $\mathbb{R}$ such that
    $\Phi(x)=\frac{1}{2}\sum_{i,j\in\mathcal{V}}W_{ij}x_{i}x_{j}+\sum_{i\in\mathcal{V}}h_{i}x_{i}+C\,,$ (5)
    for every strategy profile $x$ in $\mathcal{X}$;
  2. (iii)
    there exists a globally BR-stable set $\overline{\mathcal{N}}$ such that
    $\operatorname*{argmax}_{x\in\mathcal{X}}\Phi(x)\subseteq\overline{\mathcal{N}}\subseteq\mathcal{N}\,.$
Proof.
(i) Consider two strategy profiles $x$ and $y$ in $\mathcal{X}$ such that $x_{-i}=y_{-i}$ and $y_{i}=1=-x_{i}$ for some player $i$ in $\mathcal{V}$. Then, from (1) we have that
$u_{i}(y)-u_{i}(x)=2\sum_{k\in\mathcal{V}}W_{ik}x_{k}+2h_{i}\,.$ (6)
If $\mathcal{G}$ is undirected, then (5) and (6) imply that
$\begin{array}[]{rcl}\Phi(y)-\Phi(x)&=&\displaystyle\sum_{k\in\mathcal{V}}W_{ik}x_{k}+\sum_{k\in\mathcal{V}}W_{ki}x_{k}+2h_{i}\\[15.0pt] &=&\displaystyle 2\sum_{k\in\mathcal{V}}W_{ik}x_{k}+2h_{i}\\[15.0pt] &=&u_{i}(y)-u_{i}(x)\,,\end{array}$ (7)
thus proving that the SNC game with binary actions on $\mathcal{G}$ with external field $h$ is an exact potential game.
On the other hand, if $\mathcal{G}$ is not undirected, let $i\neq j$ in $\mathcal{V}$ be such that
$W_{ij}<W_{ji}\,,$ (8)
and consider a configuration $x$ such that $x_{i}=x_{j}=-1$. Let $y$, $w$, and $z$ in $\mathcal{X}$ be the configuration such that, respectively: $y_{i}=1$ and $y_{-i}=x_{-i}$; $w_{j}=1$ and $w_{-j}=x_{-j}$; $z_{-j}=y_{-j}$ and $z_{j}=1$. Observe that
$\begin{array}[]{rcl}u_{i}(z)-u_{i}(w)&\!\!\!=\!\!\!&\displaystyle 2\sum_{k\in\mathcal{V}}W_{ik}w_{k}+2h_{i}\\[15.0pt] &\!\!\!=\!\!\!&\displaystyle 2\sum_{k\in\mathcal{V}}W_{ik}x_{k}+2h_{i}+2\sum_{k\in\mathcal{V}}W_{ik}(w_{k}-x_{k})\\[15.0pt] &\!\!\!=\!\!\!&u_{i}(y)-u_{i}(x)+4W_{ij}\,,\end{array}$ (9)
where the last identity follows from the fact that $w_{-j}=x_{-j}$ and $w_{i}=1=-x_{j}$. Similarly, we have that
$\begin{array}[]{rcl}u_{j}(z)-u_{j}(y)&\!\!\!=\!\!\!&2\displaystyle\sum_{k\in\mathcal{V}}W_{jk}y_{k}+2h_{j}\\[15.0pt] &\!\!\!=\!\!\!&2\displaystyle\sum_{k\in\mathcal{V}}W_{jk}x_{k}+2h_{j}+2\sum_{k\in\mathcal{V}}W_{jk}(y_{k}-x_{k})\\[15.0pt] &\!\!\!=\!\!\!&u_{j}(w)-u_{j}(x)+4W_{ji}\,,\end{array}$ (10)
where the last identity follows from the fact that $y_{-i}=x_{-i}$ and $y_{i}=1=-x_{i}$. It the follows from (8), (9), and (10) that
$\begin{array}[]{rcl}0&<&4(W_{ji}-W_{ij})\\[5.0pt] &=&u_{i}(y)-u_{i}(x)+u_{j}(x)-u_{j}(w)\\[5.0pt] &&+u_{i}(w)-u_{i}(z)+u_{j}(z)-u_{j}(y)\,.\end{array}$ (11)
It then follows from [45, Corollary 2.9] that the SNC game with binary actions on $\mathcal{G}$ with external field $h$ is not an exact potential game.
(ii) If $\mathcal{G}$ is undirected, then (7) implies that $\Phi$ defined in (5) is an exact potential function for the SNC game with binary actions on $\mathcal{G}$ with external field $h$. The claim then follows from [45, Lemma 2.7].
(iii) The claim is a direct consequence of [46, Lemma 2]. ∎
III-B SNC games on unsigned networks
A finite strategic game where all players have binary action set $\mathcal{A}=\{\pm 1\}$ is super-modular[17, 18, 19, 20] if it satisfies the increasing difference property, i.e., if
$u_{i}(+1,x_{-i})-u_{i}(-1,x_{-i})\leq u_{i}(+1,y_{-i})-u_{i}(+1,y_{-i})\,,$ (12)
for every two strategy profiles $x$ and $y$ in $\mathcal{X}$ such that $x\leq y$. For super-modular games, Nash equilibria are guaranteed to exist and they form a lattice in the strategy profile space. The following result states that a SNC game is super-modular if and only if it is a network coordination game and that, in this case, the set of its Nash equilibria is non-empty and globally BR-reachable.
Proposition 2.
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$. Then,
  1. (i)
    the game is super-modular if and only if $\mathcal{G}$ is unsigned;
  2. (ii)
    if $\mathcal{G}$ is unsigned, then the set of Nash equilibria $\mathcal{N}$ is globally BR-reachable.
Proof:
(i) For every strategy profile $x$ in $\mathcal{X}$ and player $i$ in $\mathcal{V}$, let
$\delta_{i}(x)=u_{i}(+1,x_{-i})-u_{i}(+1,x_{-i})=2h_{i}+2\sum_{j\in\mathcal{V}}W_{ij}x_{j}\,.$
Then, for every two strategy profiles $x$ and $y$ in $\mathcal{X}$, we have
$\delta_{i}(y)-\delta_{i}(x)=\sum_{j\in\mathcal{V}}W_{ij}(y_{j}-x_{j})\,,$ (13)
for every $i$ in $\mathcal{V}$. If $W_{ij}\geq 0$ for every $i$ and $j$ in $\mathcal{V}$, then (13) implies that
$\delta_{i}(y)-\delta_{i}(x)=\sum_{j\in\mathcal{V}}W_{ij}(y_{j}-x_{j})\geq 0\,,$
whenever $x\leq y$, so that the increasing difference property (12) holds true, hence SNC game is super-modular.
Conversely, if $W_{ij}<0$ for some $i$ and $j$ in $\mathcal{V}$, then let $x$ and $y$ in $\mathcal{X}$ be two strategy profiles such that $x_{j}=-1$, $y_{j}=+1$, and $x_{-j}=y_{-j}$, so that $x\leq y$. Then, (13) implies that
$\delta_{i}(y)-\delta_{i}(x)=2W_{ij}<0\,,$
so that the increasing difference property (12) does not hold true, hence the SNC game is not super-modular.
(ii) This is a direct consequence of [40, Proposition 3(v)], which asserts that the set of Nash equilibria is globally I-stable (i.e., globally I-reachable and I-invariant) and the fact that I-reachability implies BR-reachability (c.f. Remark 2). ∎
Remark 3.
Proposition 2 ensures that every network coordination game, i.e., every SNC game with binary actions on an unsigned network $\mathcal{G}$, admits Nash equilibria and that the set $\mathcal{N}$ of Nash equilibria is globally reachable. Notice that, $\mathcal{N}$ may not be invariant, as Example 4 illustrates.
III-C Signed network games on structurally balanced networks
The results in the previous subsection can be extended by introducing the notion of structural balance [27, 28]. This is a property of signed networks that corresponds to the possibility of exactly partitioning the signed graph into two adversary sub-communities, such that all links within each sub-community have positive weight, whereas all links between nodes of different communities have negative weights. When the graph is structurally balanced, the SNC game can be transformed through a change of variables into a network coordination game. Consensus strategy profiles in the network coordination game corresponds to strategy profiles that take opposite sign in the two communities of the structurally balanced graph.
First, to every $\sigma$ in $\mathcal{X}=\{\pm 1\}^{\mathcal{V}}$, we can associate a diagonal matrix $[\sigma]$ in $\mathbb{R}^{\mathcal{V}\times\mathcal{V}}$ with diagonal entries
$[\sigma]_{ii}=\sigma_{i}\,,\qquad\forall i\in\mathcal{V}\,.$
Such matrices identify linear operators in $\mathbb{R}^{n}$ that are referred to as gauge transformations in some of the literature [39].
Definition 2.
For a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$, and a vector $h$ in $\mathbb{R}^{\mathcal{V}}$, consider the SNC game with binary actions on $\mathcal{G}$ with external field $h$. Given $\sigma$ in $\mathcal{X}$, let the $[\sigma]$-transformed network be
$\mathcal{G}^{[\sigma]}=(\mathcal{V},\mathcal{E},W^{[\sigma]})\,,\qquad W^{[\sigma]}=[\sigma]W[\sigma]\,,$ (14)
and the $[\sigma]$-transformed external field
$h^{[\sigma]}=[\sigma]h\,.$ (15)
Then, the $[\sigma]$-transformed game is the SNC game with binary actions on $[\sigma]$-transformed network $\mathcal{G}^{[\sigma]}$ with $[\sigma]$-transformed external field $h^{[\sigma]}$.
Observe that the utility function of every player $i$ in $\mathcal{V}$ in the $[\sigma]$-transformed game defined above is given by
$\begin{array}[]{rcl}\displaystyle u^{[\sigma]}_{i}(x)&=&\displaystyle x^{T}W^{[\sigma]}x+(h^{[\sigma]})^{T}x\\[5.0pt] &=&\displaystyle x^{T}[\sigma]W[\sigma]x+h^{T}[\sigma]x\\[5.0pt] &=&\displaystyle u_{i}([\sigma]x)\,,\end{array}$ (16)
for every strategy profile $x$ in $\mathcal{X}$. Equation (16) shows that, if we apply a gauge transformation $[\sigma]$ to a SNC game, we obtain a new SNC game whose utility functions coincide with the ones of the original game evaluated in the transformed strategy profiles $[\sigma]x$. This immediately leads to the following result.
Lemma 1.
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$, and let $\mathcal{N}$ be the set of its Nash equilibria. Given $\sigma$ in $\mathcal{X}$, consider the $[\sigma]$-transformed game and let $\mathcal{N}_{[\sigma]}$ be the set of its Nash equilibria. Then,
$\mathcal{N}=\{[\sigma]x^{*}:\,x^{*}\in\mathcal{N}_{[\sigma]}\}\,.$ (17)
Proof:
See Appendix A.∎
Gauge transformations are particularly useful when the graph is structurally balanced, as per the following definition.
Definition 3.
Consider a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$. Then:
  1. (i)
    a balanced partition of $\mathcal{G}$ is a binary partition
    $\mathcal{V}_{1}\cup\mathcal{V}_{2}=\mathcal{V}\,,\qquad\mathcal{V}_{1}\cap\mathcal{V}_{2}=\emptyset\,,$ (18)
    of its node set $\mathcal{V}$ such that
    $W_{ij}\geq 0\,,\qquad\forall i,j\in\mathcal{V}_{q}\,,\ q=1,2\,,$ (19)
    and
    $W_{ij}\leq 0\,,\qquad\forall i\in\mathcal{V}_{q}\,,\ j\in\mathcal{V}_{r}\,,\ q\neq r\,,\ q,r\in\{1,2\}\,;$ (20)
  2. (ii)
    $\mathcal{G}$ is structurally balanced if it admits a balanced partition.
Definition 3 says that a network is structurally balanced if its node set can be split into two opposite parts such that links connecting nodes in the same part all have nonnegative weight, while links connecting nodes in opposite parts all have nonpositive weight. We have the following simple result.
Lemma 2.
A network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ is structurally balanced if and only if there exists a gauge transformation $[\sigma]$ such that the $[\sigma]$-transformed network $\mathcal{G}^{[\sigma]}$ is unsigned.
Proof:
This result coincides with [39, Lemma 1]. However, since the work [39] is developed under the “digon sign-symmetry” assumption (c.f. footnote 1), and we are working in greater generality, we present a proof in Appendix B. ∎
Lemma 2 and Proposition 2 together imply the following result.
Proposition 3.
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$. If $\mathcal{G}$ is structurally balanced, then the set of Nash equilibria $\mathcal{N}$ is globally BR-reachable.
Proof:
If $\mathcal{G}$ is structurally balanced, Lemma 2 implies that there exists a gauge transformation $[\sigma]$ such that the whole node set $\mathcal{V}$ is coordinating in the $[\sigma]$-transformed network $\mathcal{G}^{[\sigma]}$. It then follows from Proposition 2(ii) that the $[\sigma]$-transformed SNC game with binary action on the network $\mathcal{G}^{[\sigma]}$ with external field $h^{[\sigma]}$ has a non-empty and globally reachable set of Nash equilibria $\mathcal{N}_{[\sigma]}$. Lemma 1 then implies that the set of Nash equilibria of the SNC game with binary actions on $\mathcal{G}$ with external field $h$ satisfies $\mathcal{N}=\{[\sigma]x^{*}:\,x^{*}\in\mathcal{N}_{[\sigma]}\}\neq\emptyset$ and is globally reachable. ∎
Example 6.
Consider the graph in Figure 4. The graph is not structurally balance. Anyway, the subset =$\mathcal{R}=\{1,\dots,4\}$ is such that the graph $\mathcal{G}_{\mathcal{R}}$ is structurally balanced. Indeed, the partition $\mathcal{R}=\mathcal{R}_{1}\cup\mathcal{R}_{2}$ with $\mathcal{R}_{1}=\{1,4\}$ and =$\mathcal{R}=\{2,3\}$ is such that $W_{\mathcal{R}_{q}\mathcal{R}_{q}}\geq 0$ for $q$ in $\{1,2\}$ and $W_{\mathcal{R}_{q}\mathcal{R}_{r}}\leq 0$ for $q\neq r$, $q$, $r$ in $\{1,2\}$. According to Proposition LABEL:pr2, there exists a gauge transformation $[\sigma]$ such that $\mathcal{R}$ is a coordinating set of $\mathcal{G}^{[\sigma]}$. Indeed, if we consider the gauge transformation $[\sigma]$ with $\sigma=[1,-1,-1,1,1,1]$, we obtain that $W^{[\sigma]}=[\sigma]W[\sigma]$ is such that
$\tilde{W}_{\mathcal{R}\mathcal{R}}$ $={\rm diag}\,\left(\begin{matrix}1\\ -1\\ -1\\ 1\end{matrix}\right)\left[\begin{matrix}0&-2&0&2\\ -2&0&1&0\\ -1&1&0&-2\\ 2&-1&0&0\end{matrix}\right]{\rm diag}\,\left(\begin{matrix}1\\ -1\\ -1\\ 1\end{matrix}\right)$
$=\left[\begin{matrix}0&2&0&2\\ 2&0&1&0\\ 1&1&0&2\\ 2&1&0&0\end{matrix}\right]\geq 0\,.$
42311-22-2-1-1561-11-1
Figure 4: Graph studied in Example 6. The subset =$\mathcal{R}=\{1,\dots,4\}$ (in gray) is such that $\mathcal{G}_{\mathcal{R}}$ is structurally balanced.
IV Main results
In this section, we present our main results.
Before proceeding, we introduce some notation that will play a crucial role in the following. For every node $i$ in $\mathcal{V}$ and subset of nodes $\mathcal{B}\subseteq\mathcal{V}$, we define
$w_{i}^{\mathcal{B}}=\sum_{j\in\mathcal{B}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|W_{ij}|}\,,$ (21)
to be its $\mathcal{B}$-out-degree. In the special case when $\mathcal{B}=\mathcal{V}$ coincides with the whole node set, we simply refer to
$w_{i}=w_{i}^{\mathcal{V}}\,,$
as the out-degree of a node $i$ in $\mathcal{V}$ and let $w=|W|\bm{1}$ be the vector of out-degrees.
We will often consider binary partitions of $\mathcal{V}$ of the type
$\mathcal{V}={\mathcal{R}}\cup{\mathcal{S}}\,,\qquad\mathcal{R}\cap\mathcal{S}=\emptyset\,,$ (22)
and identify the strategy profile space as the Cartesian product
$\mathcal{X}=\mathcal{X}_{\mathcal{R}}\times\mathcal{X}_{\mathcal{S}}\,,$
where $\mathcal{X}_{\mathcal{R}}=\mathcal{A}^{{\mathcal{R}}}$ and $\mathcal{X}_{\mathcal{S}}=\mathcal{A}^{{\mathcal{S}}}$. Correspondingly, we decompose every strategy profile $x$ in $\mathcal{X}$ as
$x=(x_{\mathcal{R}},x_{\mathcal{S}})\,,$
where $x_{\mathcal{R}}$ in $\mathcal{X}_{\mathcal{R}}$ is the strategy profile of the players in $\mathcal{R}$ and $x_{\mathcal{S}}$ in $\mathcal{X}_{\mathcal{S}}$ is the strategy profile of the players in $\mathcal{S}$.
For every strategy profile $z$ in $\mathcal{X}_{\mathcal{S}}$, we consider the SNC game with binary actions on the subnetwork $\mathcal{G}_{\mathcal{R}}$ with external field
$h^{(z)}=h_{\mathcal{R}}+W_{\mathcal{R}\mathcal{S}}z\in\mathbb{R}^{\mathcal{R}}\,,$
i.e., the game with player set ${\mathcal{R}}$, action set $\mathcal{A}=\{\pm 1\}$, and utility functions
$u_{i}^{(z)}(y)=u_{i}(y,z)=y_{i}\sum_{j\in{\mathcal{R}}}W_{ij}y_{j}+y_{i}\sum_{j\in\mathcal{S}}W_{ij}z_{j}+y_{i}h_{i}\,,$ (23)
for every player $i$ in ${\mathcal{R}}$ and strategy profile $y$ in $\mathcal{X}_{\mathcal{R}}$. We shall refer to this game as the $\mathcal{R}$-restricted SNC game with strategy profile of players in $\mathcal{S}$ frozen to $z$.
Analogously, for a given strategy profile $y$ in $\mathcal{X}_{\mathcal{R}}$, the $\mathcal{S}$-restricted game with strategy profile of players in $\mathcal{R}$ frozen to $y$ refers to the game with player set $\mathcal{S}$, action set $\mathcal{A}=\{\pm 1\}$, and utility functions
$u_{i}^{(y)}(z)=u_{i}(y,z)=z_{i}\sum_{j\in\mathcal{S}}W_{ij}z_{j}+z_{i}\sum_{j\in{\mathcal{R}}}W_{ij}y_{j}+z_{i}h_{i}\,,$ (24)
for every player $i$ in $\mathcal{S}$ and strategy profile $z$ in $\mathcal{X}_{\mathcal{S}}$. This can be interpreted as the SNC game with binary actions on the subnetwork $\mathcal{G}_{\mathcal{S}}$ with external field
$h^{(y)}=h_{\mathcal{S}}+W_{\mathcal{S}\mathcal{R}}y\in\mathbb{R}^{\mathcal{R}}\,.$
We denote by $\mathcal{N}_{\mathcal{R}}^{(z)}$ the set of Nash equilibria of the $\mathcal{R}$-restricted SNC game with strategy profile of players in $\mathcal{S}$ frozen to $z$ and by $\mathcal{N}_{\mathcal{S}}^{(y)}$ the set of Nash equilibria of the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to $y$.
IV-A Existence of Nash equilibria
In this subsection, we investigate the existence of Nash equilibria of SNC games with binary actions.
First, we consider a binary partition of the set of players as in (22), such that the subnetwork $\mathcal{G}_{\mathcal{R}}$ is unsigned, and we look for Nash equilibria of the SNC game whose projection on the coordinating set $\mathcal{R}$ is a consensus strategy profile, i.e.,
$x^{*}=(a\bm{1},z^{*})\,,$
for some $a=\pm 1$. Our first result, stated below, guarantees the existence of such Nash equilibria in terms of two properties of the $\mathcal{R}$- and $\mathcal{S}$-restricted SNC games, respcetively: the first one, cohesiveness, limits the influence that players in $\mathcal{S}$ can have on the players in $\mathcal{R}$, the second one ensures that the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to a consensus admits a Nash equilibrium.
Proposition 4.
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$, and a binary partition (22) such that $\mathcal{G}_{\mathcal{R}}$ is unsigned. Assume that there exists an action $a$ in {±$\{\pm 1\}$ such that
$w^{{\mathcal{R}}}_{i}+ah_{i}\geq w^{{\mathcal{S}}}_{i}\,,\qquad\forall i\in\mathcal{R}\,,$ (25)
and that
$\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}\neq\emptyset\,.$ (26)
Then, there exists a Nash equilibrium $x^{*}$ in $\mathcal{N}$ such that
$x^{*}_{\mathcal{R}}=a\bm{1}\,.$ (27)
Proof:
Let $x^{*}$ in $\mathcal{X}$ be a strategy profile such that $x^{*}_{\mathcal{R}}=a\bm{1}$ and $x^{*}_{\mathcal{S}}=z^{*}$ in $\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$ is a Nash equilibrium of the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{S}$ frozen to $a\bm{1}$. We will now prove that $x^{*}$ is a Nash equilibrium of the SNC game. Since $z^{*}\in\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$, what we are left to show is that $a\in\mathcal{B}_{i}(x_{-i}^{*})$ for every player $i$ in $\mathcal{R}$. Since $W_{ij}\geq 0$ for every player $i$ and $j$ in $\mathcal{R}$, from (23) this is equivalent to
$a\left(aw_{i}^{{\mathcal{R}}}+h_{i}+(W_{{\mathcal{R}}{\mathcal{S}}}z^{*})_{i}\right)\geq-a\left(aw_{i}^{{\mathcal{R}}}+h_{i}+(W_{{\mathcal{R}}{\mathcal{S}}}z^{*})_{i}\right)\,,$
or, equivalently, that
$w_{i}^{{\mathcal{R}}}+ah_{i}\geq-a(W_{{\mathcal{R}}{\mathcal{S}}}z^{*})_{i}\,,\qquad\forall i\in\mathcal{R}\,.$ (28)
Since
$w^{{\mathcal{S}}}_{i}\geq-a(W_{{\mathcal{R}}{\mathcal{S}}}z^{*})_{i}\,,\qquad\forall i\in\mathcal{R}\,,$
(28) follows from (25). This completes the proof. ∎
Remark 4.
In the special case when $h=0$, i.e., when all players are unbiased, using the identity $w_{i}^{{\mathcal{S}}}=w_{i}-w_{i}^{{\mathcal{R}}}$, we can rewrite condition (25) in Proposition 4 as
$w_{i}^{{\mathcal{R}}}\geq 1/2w_{i}\,,\qquad\forall i\in{\mathcal{R}}\,.$
This says that every node in ${\mathcal{R}}$ has at least half of the weight of its out-links towards other nodes in ${\mathcal{R}}$. In the terminology introduced in [9], this says that ${\mathcal{R}}$ is a $1/2$-cohesive subset of $\mathcal{V}$ relative to the unsigned graph $(\mathcal{V},\mathcal{E},|W|)$.
Remark 5.
Notice that the two networks of Examples 2 and 3, for which Nash equilibria did not exist, were missing just one of the two assumptions in Proposition 4. Precisely, condition (25) was not satisfied by the network in Figure 2 (a), while condition (26) was not satisfied by the network in Figure 2 (b). We also observe that global reachability is not guaranteed by conditions (25) and (26) alone, as already noticed in Example 5.
We now present the following result extending Proposition 4 to cases when $\mathcal{G}_{\mathcal{R}}$ is structurally balanced rather than unsigned. In this case, provided that $\mathcal{R}$ remains cohesive, we can determine sufficient conditions for the existence of Nash equilibria $x^{*}$ of the SNC game whose restriction $\tau=x^{*}_{\mathcal{R}}$ to the set $\mathcal{R}$ is such that the $[\tau]$-transformed subnetwork $\mathcal{G}_{\mathcal{R}}^{[\tau]}$ is unsigned.
Theorem 1.
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$. Assume that there exists a binary partition as in (22) such that $\mathcal{G}_{\mathcal{R}}$ is structurally balanced and let $\tau$ in $\mathcal{X}_{\mathcal{R}}$ be such that
$\tau_{i}W_{ij}\tau_{j}\geq 0\,,\qquad\forall i,j\in\mathcal{R}\,.$ (29)
If
$w^{{\mathcal{R}}}_{i}+\tau_{i}h_{i}\geq w^{{\mathcal{S}}}_{i}\,,\qquad\forall i\in\mathcal{R}\,,$ (30)
and
$\mathcal{N}_{{\mathcal{S}}}^{(\tau)}\neq\emptyset\,,$ (31)
then, there exists a Nash equilibrium $x^{*}$ in $\mathcal{N}$ such that
$x_{\mathcal{R}}^{*}=\tau\,.$ (32)
Proof:
Let $\sigma$ in $\mathcal{X}$ be such that $\sigma_{\mathcal{R}}=\tau$ and $\sigma_{\mathcal{S}}=\bm{1}$. Consider the associated gauge transformation $[\sigma]$ and the SNC game with binary actions on the $[\sigma]$-transformed network $\mathcal{G}^{[\sigma]}$ with external field $h^{[\sigma]}$. Observe that condition (29) ensures that $\mathcal{G}^{[\sigma]}_{\mathcal{R}}$ is unsigned. Moreover, assumption (30) implies that
$w_{i}^{\mathcal{R}}+h^{[\sigma]}_{i}\geq w_{i}^{\mathcal{S}},\qquad\forall i\in\mathcal{R}\,,$ (33)
where $w_{i}^{\mathcal{R}}$ and $w_{i}^{\mathcal{S}}$ denote, respectively, the $\mathcal{R}$- and $\mathcal{S}$-out-degrees of node $i$ (as defined in (21)) that are the same in both the original network $\mathcal{G}$ and the $[\sigma]$-transformed one $\mathcal{G}^{[\sigma]}$, since gauge transformations do not alter degrees. Furthermore, it follows from equation (16) that
$u_{i}^{[\sigma]}(\bm{1},z)=u_{i}([\sigma](\bm{1},z))=u_{i}(\tau,z)\,,\qquad\forall i\in\mathcal{S}\,,$
which, together with assumption (31), implies that
$(\mathcal{N}_{[\sigma]})_{{\mathcal{S}}}^{(\tau)}=\mathcal{N}_{{\mathcal{S}}}^{(\tau)}\neq\emptyset\,.$
Therefore, the $[\sigma]$-transformed SNC game satisfies all the assumptions of Proposition 4, hence it admits a Nash equilibrium $\tilde{x}^{*}$ in $\mathcal{N}_{[\sigma]}$ such that $\tilde{x}^{*}_{\mathcal{R}}=\bm{1}$. Lemma 1 then implies that $x^{*}=[\sigma]\tilde{x}^{*}\in\mathcal{N}$ is a Nash equilibrium for the SNC game with binary actions on $\mathcal{G}$ with external field $h$. Finally, observe that
$x^{*}_{\mathcal{R}}=([\sigma]\tilde{x}^{*})_{\mathcal{R}}=[\tau]\bm{1}=\tau\,,$
so that (32) holds true, thus completing the proof. ∎
Observe that assumption (31) in Theorem 1, namely the fact that the set $\mathcal{N}_{{\mathcal{S}}}^{(\tau)}$ of Nash equilibria of the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to $\tau$, is automatically verified when the subnetwork $\mathcal{G}_{\mathcal{S}}$ is either undirected or structurally balanced itself. This leads to the following corollaries.
Corollary 1.
Consider a SNC game with binary actions on $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ and a binary partition (22) where both $\mathcal{G}_{\mathcal{R}}$ and $\mathcal{G}_{\mathcal{S}}$ are structurally balanced. If there exists $\tau$ in $\mathcal{X}_{\mathcal{R}}$ such that assumption (30) holds true, then there exists a Nash equilibrium $x^{*}\in\mathcal{N}$ satisfying equation (32).
Proof:
If the subnetwork $\mathcal{G}_{\mathcal{S}}$ is structurally balanced, then, for every $\tau$ in $\mathcal{X}_{\mathcal{R}}$, Proposition 3 applied to the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to $\tau$, implies that assumption (31) holds true. The claim then follows from Theorem 1. ∎
Corollary 2.
Consider a SNC game with binary actions on $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ and a binary partition (22) where $\mathcal{G}_{\mathcal{R}}$ is structurally balanced and $\mathcal{G}_{\mathcal{S}}$ is undirected. If there exists $\tau$ in $\mathcal{X}_{\mathcal{R}}$ such that condition (30) holds true, then there exists a Nash equilibrium $x^{*}\in\mathcal{N}$ satisfying equation (32).
Proof:
If the subnetwork $\mathcal{G}_{\mathcal{S}}$ is undirected, then, for every $\tau$ in $\mathcal{X}_{\mathcal{R}}$, Proposition 1 applied to the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to $\tau$ implies that assumption (31) holds true. The claim then follows from Theorem 1. ∎
Remark 6.
Corollaries 1 and 2 significantly generalize previous results where the existence of (pure strategy) Nash equilibria was proved only for network coordination or network anti-coordination games over undirected graphs [22]. Moreover, Corollary 2 applies to the mixed network coordination/anti-coordination games studied in [1, 2].
Example 1 (cont’d).
Consider the graph in Figure 1 and let $h=0$. Observe that, for all $i$ in $\mathcal{R}$, it holds that $w_{i}^{\mathcal{R}}\geq w_{i}^{\mathcal{S}}$. Since the subnetwork $\mathcal{G}_{\mathcal{R}}$ is unsigned and the subsetwork $\mathcal{G}_{\mathcal{S}}$ is undirected, Corollary 2 implies the existence of two Nash equilibria $x^{*}$ and $x^{**}$ in $\mathcal{N}$ such that $x^{*}_{\mathcal{R}}=\bm{1}=-x^{**}_{\mathcal{R}}$.
IV-B Stability of Nash equilibria
We now present results on the global stability of Nash equilibria of SNC games. Before proceeding, it is convenient to reconsider the SNC game in Example 5, whose set of Nash equilibria was shown to be not globally BR-reachable. A closer look at this example suggests that this is a direct consequence of the topological structure of the underlying network displayed in Figure 3. This network contains two components $\{1,2,3\}$ and $\{5,6,7\}$, each of which without any out-link towards other nodes in the graph. This decomposition of the graph is what prevents the coordinating players to reach a consensus starting from a polarized initial condition. The observation above motivates the following definition that introduces a property of the graph guaranteeing that similar decompositions are not possible. The proposed definition also accounts for the external field $h$. It will be at the basis of the results presented in this subsection.
Definition 4.
Consider a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ and two vectors $h^{-}$ and $h^{+}$ in $\mathbb{R}^{\mathcal{V}}$ such that $h^{-}\leq h^{+}$. We say that $\mathcal{G}$ is $(h^{-},h^{+})$-indecomposable if for every binary partition
$\mathcal{V}=\mathcal{V}^{-}\cup\mathcal{V}^{+}\,,\qquad\mathcal{V}^{-}\cap\mathcal{V}^{+}=\emptyset\,,\qquad\mathcal{V}^{-}\neq\emptyset\neq\mathcal{V}^{+}\,,$ (34)
there exists a node $i$ in $\mathcal{V}$ such that either
$i\in\mathcal{V}^{+}\text{ and }w_{i}^{\mathcal{V}^{+}}+h_{i}^{+}<w_{i}^{\mathcal{V}^{-}}$ (35)
or
$i\in\mathcal{V}^{-}\text{ and }w_{i}^{\mathcal{V}^{-}}-h_{i}^{-}<w_{i}^{\mathcal{V}^{+}}\,.$ (36)
The following example illustrates the notion of indecomposability introduced in Definition 4 above.
4231122111
Figure 5: Graph considered in Example 7.
Example 7.
The network $\mathcal{G}$ in Figure 5 is $(h^{-},h^{+})$-indecomposable for $h^{+}=(3,2,0,2)=-h^{-}$. To see this, first notice that, for any partitiion where $|\mathcal{V}^{+}|=|\mathcal{V}^{-}|=2$, node $3$ satifies either (35) or (36) as a direct consequence of the facts that its out-degree is $w_{3}=3$ and $h_{3}^{-}=h_{3}^{+}=0$. Suppose instead that $|\mathcal{V}^{-}|=1$. A direct check shows that the node in $\mathcal{V}^{-}$ satisfies (36). The case when $|\mathcal{V}^{+}|=1$ is identical.
The network $\mathcal{G}$ in Figure 5 is not $(h^{-},h^{+})$-indecomposable for $h^{+}=\bm{1}$ and $h^{-}=-h^{+}$. Indeed, if we take the partition $\mathcal{V}^{+}=\{1,4\}$ and $\mathcal{V}^{-}=\{2,3\}$, we have that $w_{1}^{\mathcal{V}^{+}}+h_{1}^{+}=3\geq 2=w_{1}^{\mathcal{V}^{-}}$, $w_{4}^{\mathcal{V}^{+}}+h_{4}^{+}=3\geq 1=w_{4}^{\mathcal{V}^{-}}$, $w_{2}^{\mathcal{V}^{+}}-h_{2}^{-}=2\geq 1=w_{2}^{\mathcal{V}^{+}}$ and $w_{3}^{\mathcal{V}^{-}}-h_{3}^{-}=2\geq 2=w_{3}^{\mathcal{V}^{+}}$. Then, $\nexists i$ in $\mathcal{V}^{+}$ or $\mathcal{V}^{-}$ such that either (35) or (36) is violated.
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$ such that
$h^{-}_{i}\leq h_{i}\leq h_{i}^{+}\,,\qquad\forall i\in\mathcal{V}\,.$ (37)
Given a binary partition as in (34), consider the strategy profile $x$ in $\mathcal{X}$ with
$x_{\mathcal{V}^{+}}=\bm{1}\,,\qquad x_{\mathcal{V}^{-}}=-\bm{1}\,.$
[…]
We will make use of a result in [40] that ensures, for network coordination games on $(h^{-},h^{+})$-indecomposable unsigned networks, the existence of a BR-path from every strategy profile $x$ to a consensus strategy profile that is independent from the specific choice of the vector $h$ satisfying (37). Precisely, [40, Theorem 4(i)] implies the following.
Lemma 3.
Consider an unsigned network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ and two vectors $h^{-}$, $h^{+}$ in $\mathbb{R}^{\mathcal{V}}$ for which $\mathcal{G}$ is $(h^{-},h^{+})$-indecomposable. Then, for every strategy profile $x^{(0)}$ in $\mathcal{X}$, there exists an $l$-tuple of strategy profiles $(x^{(1)},\ldots,x^{(l)})$, with $1\leq l\leq n$, such that $x^{(l)}\in\{\pm\bm{1}\}$ is a consensus profile, and $(x^{(0)},x^{(1)},\ldots,x^{(l)})$ is a BR-path for every SNC game with binary actions on $\mathcal{G}$ with external field $h$ satisfying (37).
We can now get the following result.
Proposition 5.
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$ and a binary partition (22) such that $\mathcal{G}_{\mathcal{R}}$ is unsigned. Let $h^{-}$ and $h^{+}$ in $\mathbb{R}^{\mathcal{R}}$ be the vectors with entries
$h^{+}_{i}=h_{i}+w_{i}^{{\mathcal{S}}}\,,\qquad h^{-}_{i}=h_{i}-w_{i}^{{\mathcal{S}}}\,,\qquad\forall i\in\mathcal{R}\,.$ (38)
Assume that $\mathcal{G}_{\mathcal{R}}$ is $(h^{-},h^{+})$-indecomposable and
$w^{{\mathcal{R}}}_{i}-|h_{i}|>w^{{\mathcal{S}}}_{i}\,,\qquad\forall i\in\mathcal{R}\,.$ (39)
Then:
  1. (i)
    if, for $y=\pm\bm{1}$ in $\mathcal{X}_{\mathcal{R}}$, the set $\mathcal{N}_{{\mathcal{S}}}^{(y)}$ is globally BR-reachable for the $\mathcal{S}$-restricted SNC game with action profile of players in $\mathcal{R}$ frozen to $y$, then, the subset of Nash equilibria
    $\tilde{\mathcal{N}}=\{x^{*}\in\mathcal{N}:\,x^{*}_{\mathcal{R}}\in\{\pm\bm{1}\}\}\,,$ (40)
    is non-empty and globally BR-reachable.
Moreover:
  1. (ii)
    if, for $y=\pm\bm{1}$ in $\mathcal{X}_{\mathcal{R}}$, there exists a non-empty subset $\bar{\mathcal{N}}_{{\mathcal{S}}}^{(y)}\subseteq\mathcal{N}_{{\mathcal{S}}}^{(y)}$ that is globally BR-stable for the $\mathcal{S}$-restricted SNC game with action profile of players in $\mathcal{R}$ frozen to $y$, then there exists a non-empty globally BR-stable subset $\overline{\mathcal{N}}\subseteq\tilde{\mathcal{N}}$.
Proof:
Fix an arbitrary strategy profile $x$ in $\mathcal{X}$ and let $z=x_{\mathcal{S}}$. On the one hand, since the subnetwork $\mathcal{G}_{\mathcal{R}}$ is unsigned and $(h^{-},h^{+})$-indecomposable, Lemma 3 implies that the set of consensus strategy profiles {±$\{\pm\bm{1}\}\subseteq\mathcal{X}_{\mathcal{R}}$ is globally BR-reachable for the $\mathcal{R}$-restricted network coordination game with action profile of players in $\mathcal{S}$ frozen to $z$, so that there exists a length-$l$ BR-path $((y^{(0)},z),(y^{(1)},z),\ldots,(y^{(l)},z))$ with $y^{(0)}=x_{\mathcal{R}}$ and $y^{(l)}=a\bm{1}$, for some $a$ in {±$\{\pm 1\}$. On the other hand, since the set $\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$ is globally BR-reachable for the $\mathcal{S}$-restricted SNC game with action profile of players in $\mathcal{R}$ frozen to $y=a\bm{1}$, there exists a length-$m$ BR-path $((a\bm{1},z^{(0)}),(a\bm{1},z^{(1)}),\ldots,(a\bm{1},z^{(m)}))$ with $z^{(0)}=z$ and $z^{(m)}=z^{*}\in\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$.
Observe that the strategy profile $x^{*}$ in $\mathcal{X}$ with $x^{*}_{\mathcal{R}}=a\bm{1}$ and $x^{*}_{\mathcal{S}}=z^{*}$ is a Nash equilibrium for the original SNC game with binary actions on $\mathcal{G}$ with external field $h$, as every player $i$ in $\mathcal{R}$ is playing best response thanks to (39), while every player $j$ in $\mathcal{S}$ is playing best response since $z^{*}\in\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$. We have thus found a length-$(l+m)$ BR-path $((y,z),(y^{(1)},z),\ldots,(a\bm{1},z),(a\bm{1},z^{(1)}),\ldots,(a\bm{1},z^{*}))$ from $x$ to $x^{*}$ in $\mathcal{N}$ with $x^{*}_{\mathcal{R}}=a\bm{1}$. The arbitrariness of initial strategy profile $x$ in $\mathcal{X}$ implies that the set $\tilde{\mathcal{N}}$ defined in (40), i.e., the subset of Nash equilibria in which players in $\mathcal{R}$ are at consensus is globally BR-reachable, thus proving point (i) of the claim.
To prove point (ii) of the claim, observe that, if there exists a non-empty subset $\bar{\mathcal{N}}_{{\mathcal{S}}}^{(a\bm{1})}\subseteq\mathcal{N}_{{\mathcal{S}}}^{(a\bm{1})}$ that is globally BR-stable for the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to $y=a\bm{1}$, then the BR-path above can be constructed leading to $x^{*}$ such that $x^{*}_{\mathcal{S}}\in\bar{\mathcal{N}}_{{\mathcal{S}}}^{(a\bm{1})}$. Now, notice that assumption (39) implies that, in the strategy profile $x^{*}$ defined above, every player $i$ in ${\mathcal{R}}$ is playing a strict best response $\mathcal{B}_{i}(x^{*}_{-i})=\{a\}$. This, combined with the BR-invariance of $\overline{\mathcal{N}}_{\mathcal{S}}^{(a\bm{1})}$ for the $\mathcal{S}$-restricted SNC game withstrategy profile of players in $\mathcal{R}$ frozen to $y=a\bm{1}$, implies that
$\overline{\mathcal{N}}=\{x^{*}\in\mathcal{X}:\,x^{*}_{\mathcal{R}}\in\{\pm\bm{1}\}\,,\,x^{*}_{\mathcal{S}}\in\bar{\mathcal{N}}_{{\mathcal{S}}}^{(x^{*}_{\mathcal{R}})}\}\,,$
is a non-empty, globally BR-stable subset of Nash equilibria, thus proving point (ii) of the claim. ∎
Remark 7.
In the special case of a network coordination game, i.e., when the network $\mathcal{G}$ is unsigned, Proposition 5 provides sufficient conditions for consensus strategy profiles ±$\pm\bm{1}$ to be Nash equilibria and form a globally BR-reachable set. In this setting, the assumptions reduce to the following two conditions: (a) that $\mathcal{G}_{\mathcal{R}}$ is $(h^{-},h^{+})$-indecomposable; and (b) that $w_{i}>|h_{i}|$ for every $i$ in $\mathcal{V}$. Condition (a) ensures that no coexistent Nash equilibrium exists, while condition (b) ensures that both consensus strategy profiles ±$\pm\bm{1}$ are strict Nash equilibria. We notice that in this case conditions (a) and (b) are not just sufficient but also necessary for the set {±$\{\pm\bm{1}\}$ of consensus strategy profiles to be globally BR-stable.
We now present the following result extending Proposition 5 to cases when $\mathcal{G}_{\mathcal{R}}$ is structurally balanced rather than unsigned.
Theorem 2.
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$. Consider a binary partition as in (22), such that $\mathcal{G}_{\mathcal{R}}$ is structurally balanced and let $\tau$ in $\mathcal{X}_{\mathcal{R}}$ be such that $\mathcal{G}_{\mathcal{R}}^{[\tau]}$ is unsigned. Let $h^{-}$ and $h^{+}$ in $\mathbb{R}^{\mathcal{R}}$ be the vectors with entries
$h^{+}_{i}=\tau_{i}h_{i}+w_{i}^{{\mathcal{S}}}\,,\qquad h^{-}_{i}=\tau_{i}h_{i}-w_{i}^{{\mathcal{S}}}\,,\qquad\forall i\in\mathcal{R}\,.$ (41)
Assume that $\mathcal{G}_{\mathcal{R}}^{[\tau]}$ is $(h^{-},h^{+})$-indecomposable and that (39) holds true. Then:
  1. (i)
    if, for $y=\pm\tau$ in $\mathcal{X}_{\mathcal{R}}$, the set $\mathcal{N}_{{\mathcal{S}}}^{(y)}$ is globally BR-reachable for the $\mathcal{S}$-restricted SNC game with the strategy profile in $\mathcal{R}$ frozen to $y$, then, the subset of Nash equilibria
    $\overline{\mathcal{N}}=\{x^{*}\in\mathcal{N}:x^{*}_{\mathcal{R}}\in\{\pm\tau\}\}\,,$ (42)
    is globally BR-reachable.
Moreover:
  1. (ii)
    if, for $y=\pm\tau$ in $\mathcal{X}_{\mathcal{R}}$, there exists a non-empty subset $\bar{\mathcal{N}}_{{\mathcal{S}}}^{(y)}\subseteq\mathcal{N}_{{\mathcal{S}}}^{(y)}$ that is globally BR-stable for the $\mathcal{S}$-restricted SNC game where actions of players in $\mathcal{R}$ are frozen to $y$, then there exists a globally BR-stable subset of Nash equilibria contained in $\overline{\mathcal{N}}$.
Proof:
Let $\sigma$ in $\mathcal{X}$ be such that $\sigma_{\mathcal{R}}=\tau$ and $\sigma_{\mathcal{S}}=\bm{1}$. Consider the associated gauge transformation $[\sigma]$ and the $[\sigma]$-transformed network $\mathcal{G}^{[\sigma]}$. Observe that the assumptions ensure that $(\mathcal{G}^{[\sigma]})_{\mathcal{R}}=\mathcal{G}_{\mathcal{R}}^{[\tau]}$ is unsigned, that $\mathcal{G}_{\mathcal{R}}$ is $(h^{-},h^{+})$-indecomposable and that (39) holds true. We can thus apply Proposition (5) to the $[\sigma]$-transformed SNC game, whose points (i) and (ii) imply, respectively, points (i) and (ii) of the claim. ∎
Similarly to Section IV-A, we may derive the following two corollaries from Theorem 2.
Corollary 3.
Consider a SNC game on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$. Consider a binary partition as in (22) such that $\mathcal{G}_{\mathcal{R}}$ and $\mathcal{G}_{\mathcal{S}}$ are both structurally balanced, and let $\tau$ in $\mathcal{X}_{\mathcal{R}}$ be such that $\mathcal{G}_{\mathcal{R}}^{[\tau]}$ is unsigned. Let $h^{-}$ and $h^{+}$ in $\mathbb{R}^{\mathcal{R}}$ be the vectors with entries as in (38). If $\mathcal{G}_{\mathcal{R}}$ is $(h^{-},h^{+})$-indecomposable and (39) holds true, then the subset of Nash equilibria (42) is globally BR-reachable.
Proof:
If the subnetwork $\mathcal{G}_{\mathcal{S}}$ is structurally balanced, then, for every $\tau$ in $\mathcal{X}_{\mathcal{R}}$, Proposition 3 implies that the set $\mathcal{N}_{{\mathcal{S}}}^{(y)}$ is globally BR-reachable for the $\mathcal{S}$-restricted SNC game with the strategy profile in $\mathcal{R}$ frozen to $y=\pm\tau$. The claim then follows from Theorem 2(i). ∎
4231122111652-1-12-1
Figure 6: Signed graph with coordinating set =$\mathcal{R}=\{1,\dots,4\}$ in gray (see Example 8).
Example 8.
Consider a SNC game on the graph in Figure 6 with $h=0$. Observe that =$\mathcal{R}=\{1,\dots,4\}$ and $\mathcal{S}=\mathcal{V}\setminus\mathcal{R}$ are coordinating sets with $w^{\mathcal{R}}|_{\mathcal{R}}-\lvert d\rvert_{\mathcal{R}}=(4,3,3,3)>w^{\mathcal{S}}|_{\mathcal{R}}=(2,2,0,1)$. Furthermore, the graph $\mathcal{G}_{\mathcal{R}}$ is $(h^{+},h^{-})$-indecomposable for $h^{+}=h_{\mathcal{R}}+w^{\mathcal{S}}|_{\mathcal{R}}=(2,2,0,1)$ and $h^{-}=h_{\mathcal{R}}-w^{\mathcal{S}}|_{\mathcal{R}}=-h^{+}$. This can be proved following the same reasoning as in Example 7 (notice that $\mathcal{G}_{\mathcal{R}}$ coincides with the graph in Figure 7). Then, according to Corollary 3, the set of Nash equilibria where the players in ${\mathcal{R}}$ are at consensus is globally BR-reachable. Notice that the set of Nash equilibria
$\mathcal{N}=\{(a\bm{1}_{\mathcal{R}},-a,a),(a\bm{1}_{\mathcal{R}},a,-a),a=\pm 1\}\,.$
is not globally BR-stable. Indeed, for every $x^{*}$ in $\mathcal{N}$, the best response of player $5$ is $\mathcal{B}_{5}(x_{-5}^{*})=\{\pm 1\}$, while $6$ is playing a strict best response. Therefore, there exists a best-response path (e.g., $((a\bm{1}_{\mathcal{R}},-a,a),(a\bm{1}_{\mathcal{R}},a,a))$ for $a=\pm 1$) that leaves the set of Nash equilibria.
Corollary 4.
Consider a SNC game on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$. Consider a binary partition as in (22) such that $\mathcal{G}_{\mathcal{R}}$ is structurally balanced and $\mathcal{G}_{\mathcal{S}}$ is undirected. Let $\tau$ in $\mathcal{X}_{\mathcal{R}}$ be such that $\mathcal{G}_{\mathcal{R}}^{[\tau]}$ is unsigned and let $h^{-}$ and $h^{+}$ in $\mathbb{R}^{\mathcal{R}}$ be the vectors with entries as in (38). If $\mathcal{G}_{\mathcal{R}}$ is $(h^{-},h^{+})$-indecomposable and (39) holds true, then there exists a globally BR-stable subset of Nash equilibria.
Proof:
If the subnetwork $\mathcal{G}_{\mathcal{S}}$ is undirected, then Proposition 1 implies that, for $y=\pm\tau$ in $\mathcal{X}_{\mathcal{R}}$, there exists a non-empty subset $\bar{\mathcal{N}}_{{\mathcal{S}}}^{(y)}\subseteq\mathcal{N}_{{\mathcal{S}}}^{(y)}$ that is globally BR-stable for the $\mathcal{S}$-restricted SNC game where actions of players in $\mathcal{R}$ are frozen to $y$. The claim then follows from Theorem 2(ii). ∎
423112211165-11-1-1-1
Figure 7: Graph studied in Example 9.
Example 9.
Consider the SNC game on the graph in Figure 7 with $h=(2,0,0,-1,0,0)$. Observe that =$\mathcal{R}=\{1,\dots,4\}$ is a coordinating set with $w^{\mathcal{R}}|_{\mathcal{R}}=(4,3,3,3)$ and $w^{\mathcal{S}}|_{\mathcal{R}}=(1,2,0,1)$ and $\mathcal{G}_{\mathcal{S}}$ with $\mathcal{S}=\mathcal{V}\setminus\mathcal{R}$ is undirected. It holds that
$w^{\mathcal{R}}|_{\mathcal{R}}-\lvert h\rvert_{\mathcal{R}}=(2,3,3,2)>w^{\mathcal{S}}|_{\mathcal{R}}=(1,2,0,1)\,.$
Furthermore, the graph $\mathcal{G}_{\mathcal{R}}$ is $(h^{-},h^{+})$-indecomposable for $\tilde{h}^{+}=h|_{\mathcal{R}}+w^{\mathcal{S}}|_{\mathcal{R}}=(3,2,0,0)$ and $\tilde{h}^{-}=h|_{\mathcal{R}}-w^{\mathcal{S}}|_{\mathcal{R}}=(1,-2,0,-2)$. Again, this can proved following the same reasoning as in Example 7 and 8. Then, according to Corollary 4, the set of Nash equilibria where the players in ${\mathcal{R}}$ are at consensus contains a globally BR-stable subset.
Example 10.
Consider now the SNC game on the graph in Figure 8 and let $h=0$. Analogously to Example 9, we have that =$\mathcal{R}=\{1,\dots,4\}$ is a coordinating set with $w^{\mathcal{R}}|_{\mathcal{R}}=(4,3,3,3)$, and $\mathcal{G}_{\mathcal{S}}$ with $\mathcal{S}=\mathcal{V}\setminus\mathcal{R}$ is undirected. On the other hand, in this case, we have $w^{\mathcal{S}}|_{\mathcal{R}}=(1,1,1,1)$. Since $w^{\mathcal{R}}|_{\mathcal{R}}\geq w^{\mathcal{S}}|_{\mathcal{R}}$, Corollary 2 holds and existence of a Nash equilibrium where the players in $\mathcal{R}$ are at consensus is guaranteed. However, $\mathcal{G}_{\mathcal{R}}$ is not $(h^{+},h^{-})$-indecomposable for $h^{+}=h+w^{\mathcal{S}}|_{\mathcal{R}}=(1,1,1,1)$ and $h^{-}=h-w^{\mathcal{S}}|_{\mathcal{R}}=-h^{+}$. Indecomposability is indeed violated by +=$\mathcal{R}^{+}=\{1,4\}$ and =$\mathcal{R}^{-}=\{2,3\}$. Therefore, Proposition 5 and Theorem 2 do not apply. Observe that $x^{*}=(1,-1,-1,1,-1,1,1,-1)$ is a strict Nash equilibrium of the game where players in $\mathcal{R}$ are not at consensus.
423112211156-1-1-1-178-3-3
Figure 8: Graph studied in Example 10
Example 6 (cont’d).
Consider the graph in Figure 4. As observed in Example 6 subset =$\mathcal{R}=\{1,\dots,4\}$ is such that the graph $\mathcal{G}_{\mathcal{R}}$ is structurally balanced. Indeed, if we consider the gauge transformation $[\sigma]$ with $\sigma=[1,-1,-1,1,1,1]$, we obtain that $W^{[\sigma]}=[\sigma]W[\sigma]\geq 0$. Furthermore, for $h_{i}=0$, we have $w^{\mathcal{R}}|_{\mathcal{R}}=(4,3,4,3)>w^{\mathcal{S}}|_{\mathcal{R}}=(1,1,1,1)$. Since $\mathcal{G}_{\mathcal{S}}$ is a undirected (no links), this implies that, for $\tau=(1,-1,-1,1)$, there exists a polarized Nash equilibrium, that is, a Nash equilibrium where $x_{i}^{*}=\tau_{i}$ for all $i$ in $\mathcal{R}$. For instance, the strategy profile $x^{*}=(1,-1,-1,1,-1,1)$ is a Nash equilibrium of the game.
Finally, since $\tilde{\mathcal{G}}_{\mathcal{R}}$ with $\tilde{W}_{\mathcal{R}\mathcal{R}}=\lvert W_{\mathcal{R}\mathcal{R}}\rvert$ is $(h^{+},h^{-})$-indecomposable for $h^{+}=\bm{1}$ and $h^{-}=-\bm{1}$, the set of Nash equilibria that are polarized in the structurally balanced component admits a globally BR-stable subset. We remark that these results can be obtained without knowing the partitions $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$.
V Concluding remarks
The signed network coordination games studied in this paper encompass a number of network strategic games that have appeared in the recent literature. They model the contemporaneous presence of strategic complement and strategic substitute effects in an economic multi-player model, or rather the presence of antagonistic behaviors in a social network. Such games pose challenging problems, as Nash equilibria may not exist and even when they exist the behavior of learning dynamics may be complex and sensitive to initial condition and the order of activation of the various players. In this paper, we have obtained conditions under which a subset of coordinating players is capable of forcing the convergence of best response dynamics to a Nash equilibrium that is a consensus on their part. Our results use in a novel way the concept of cohesiveness proposed in [9] and build on the super-modular properties of coordinating games. Further work includes finding efficient algorithms to verify the proposed conditions and deriving necessary conditions for existence, reachability and stability of Nash equilibria.
Acknowledgments
The authors gratefully acknowledge Prof. Claudio Altafini and Prof. Angela Fontan for insightful discussions.
Appendix A Proof of Lemma 1
First, observe that $x^{*}$ in $\mathcal{X}$ is a Nash equilibrium for the $[\sigma]$-transformed game if and only if
$x_{i}^{*}\left(\sum\nolimits_{j}W^{[\sigma]}_{ij}x^{*}_{j}+h^{[\sigma]}_{i}\right)\geq 0\,,\qquad\forall i\in\mathcal{V}\,.$
Now, notice that
$\begin{array}[]{rcl}\!\!\!\displaystyle x_{i}^{*}\left(\sum\nolimits_{j}W^{[\sigma]}_{ij}x^{*}_{j}+h^{[\sigma]}_{i}\right)&\!\!\!\!=\!\!\!\!&x_{i}^{*}\left(\displaystyle\sum\nolimits_{j}\sigma_{i}W_{ij}\sigma_{j}x^{*}_{j}+\sigma_{i}h_{i}\right)\\[10.0pt] &\!\!\!\!=\!\!\!\!&\displaystyle\sigma_{i}x_{i}^{*}\left(\sum\nolimits_{j}W_{ij}\sigma_{j}x^{*}_{j}+h_{i}\right)\\[10.0pt] &\!\!\!\!=\!\!\!\!&\displaystyle([\sigma]x^{*})_{i}\left(\sum\nolimits_{j}W_{ij}([\sigma]x^{*})_{j}+h_{i}\right).\end{array}$
Therefore, $x^{*}$ is a Nash equilibrium for the $[\sigma]$-transformed game if and only if $[\sigma]x^{*}$ is a Nash equilibrium for the SNC game with binary actions on $\mathcal{G}$ with external field $h$.
Appendix B Proof of Lemma 2
If $\mathcal{G}$ is structurally balanced, then consider a balanced partition as in (18) and let $\sigma$ in $\mathcal{X}$ have entries $\sigma_{i}=-1$ for every $i$ in $\mathcal{V}_{1}$ and $\sigma_{i}=+1$ for every $i$ in $\mathcal{V}_{2}$. It then follows from (19) that $W^{[\sigma]}_{ij}=\sigma_{i}W_{ij}\sigma_{j}=W_{ij}\geq 0$, for every $i$ and $j$ in $\mathcal{V}_{q}$, for $q=1,2$, whereas (20) implies that $W^{[\sigma]}_{ij}=\sigma_{i}W_{ij}\sigma_{j}=-W_{ij}\geq 0$, for every $i$ in $\mathcal{V}_{q}$ and $j$ in $\mathcal{V}_{r}$, for $q\neq r$, $q,r=1,2$. This shows that, if $\mathcal{G}$ is structurally balanced, then $W^{[\sigma]}$ is a nonnegative matrix, hence the whole set $\mathcal{V}$ is coordinating for the transformed network $\mathcal{G}^{[\sigma]}$.
Conversely, let $[\sigma]$ be a gauge transformation such that the whole node set $\mathcal{V}$ is coordinating for the transformed network $\mathcal{G}^{[\sigma]}$, i.e., $W^{[\sigma]}_{ij}\geq 0$ for every $i$ and $j$ in $\mathcal{V}$. Define
$\mathcal{V}_{1}=\{i\in\mathcal{V}\!:\,\sigma_{i}=-1\}\,,\qquad\mathcal{V}_{2}=\{i\in\mathcal{V}\!:\,\sigma_{i}=1\}\,.$ (43)
Then, clearly (18) holds true. Moreover, for every $i$ and $j$ in $\mathcal{V}_{q}$, for $q=1,2$, we have $W_{ij}=\sigma_{i}W^{[\sigma]}_{ij}\sigma_{j}=W_{ij}^{[\sigma]}\geq 0$, so that (19) holds true. Furthermore, for every $i$ in $\mathcal{V}_{q}$ and $j$ in $\mathcal{V}_{r}$, for $q\neq r$, we have $W_{ij}=\sigma_{i}W^{[\sigma]}_{ij}\sigma_{j}=-W^{[\sigma]}_{ij}\geq 0$, so that (20) holds true as well. Therefore, (43) determines a balanced partition of the node set $\mathcal{V}$, so that $\mathcal{G}$ is structurally balanced. ∎
[Uncaptioned image] Martina Vanelli received the B.Sc., M.S. (cum laude), and Ph.D. (with honors) degrees in Applied Mathematics from Politecnico di Torino, in 2017, 2019, and 2024, respectively. She is currently a postdoctoral fellow at the Institute for Information and Communication Technologies, Electronics and Applied Mathematics (ICTEAM), Université catholique de Louvain. From October 2018 to March 2019, she was a visiting student at Technion, Israel Institute of Technology. Her research interests include identification, analysis, and control of multi-agent systems, with application to social and economic networks, and power markets.
[Uncaptioned image] Laura Arditti graduated with honors as a PhD in Applied Mathematics at the Department of Mathematical Sciences of Politecnico di Torino in 2023. She previously received the B.Sc. in Physics Engineering in 2016 and the M.S. in Mathematical Engineering in 2018, both magna cum laude from Politecnico di Torino. She currently works in the financial sector as quantitative Financial Crime Prevention analyst. Her research focused on game theory, its relationship with graphical models, and its applications to economic and financial networks.
[Uncaptioned image] Giacomo Como is a Professor at the Department of Mathematical Sciences, Politecnico di Torino, Italy, and a Senior Lecturer at the Automatic Control Department of Lund University, Sweden. He received the B.Sc., M.S., and Ph.D. degrees in Applied Mathematics from Politecnico di Torino, in 2002, 2004, and 2008, respectively. He was a Visiting Assistant in Research at Yale University in 2006–2007 and a Postdoctoral Associate at the Laboratory for Information and Decision Systems, Massachusetts Institute of Technology, from 2008 to 2011. He currently serves as Senior Editor of the IEEE Transactions on Control of Network Systems and Associate Editor of Automatica. . He has been serving as Associate Editor of the IEEE Transactions on Network Science and Engineering and of the IEEE Transactions on Control of Network Systems and as the chair of the IEEE-CSS Technical Committee on Networks and Communications. He was the IPC chair of the IFAC Workshop NecSys’15 and a semi-plenary speaker at the International Symposium MTNS’16. He is recipient of the 2015 George S.  Axelby Outstanding Paper Award. His research interests are in dynamics, information, and control in network systems with applications to cyber-physical systems, infrastructure networks, and social and economic networks.
[Uncaptioned image] Fabio Fagnani received the Laurea degree in Mathematics from the University of Pisa and the Scuola Normale Superiore, Pisa, Italy, in 1986. He received the PhD degree in Mathematics from the University of Groningen, Groningen, The Netherlands, in 1991. From 1991 to 1998, he was an Assistant Professor of Mathematical Analysis at the Scuola Normale Superiore. In 1997, he was a Visiting Professor at the Massachusetts Institute of Technology (MIT), Cambridge, MA. Since 1998, he has been with the Politecnico of Torino, where since 2002 he has been a Full Professor of Mathematical Analysis. From 2006 to 2012, he has acted as Coordinator of the PhD program in Mathematics for Engineering Sciences at Politecnico di Torino. From June 2012 to September 2019, he served as the Head of the Department of Mathematical Sciences, Politecnico di Torino. His current research topics are on cooperative algorithms and dynamical systems over networks, inferential distributed algorithms, and opinion dynamics. He is an Associate Editor of the IEEE Transactions on Automatic Control and served in the same role for the IEEE Transactions on network Science and Engineering and of the IEEE Transactions on Control of network Systems.